OpenStudy (anonymous):

Multiply. Write the product in simplest form 3/4 (-8/25)(-5/12)

OpenStudy (anonymous):

1,10th

OpenStudy (anonymous):

1 10

OpenStudy (anonymous):

( 3 4 ( −8 25 ))( −5 12 ) = 3 4 ( −8 25 )( −5 12 ) = 3 4 ( −8 25 )( −5 12 ) = −6 25 ( −5 12 ) = 1 10 this is it step by step

OpenStudy (anonymous):

its confusing looking at it like that

Directrix (directrix):

@mathwiz916 Please come to this thread and explain your work. We cannot understand it. Thanks.

OpenStudy (jdoe0001):

probably becuause he/she simply copied and pasted it from somewhere as opposed to actually do it and the paste came out out of whack

OpenStudy (anonymous):

yes i do agree with them

OpenStudy (jdoe0001):

ahemm \(\bf \cfrac{3}{4}\left( -\cfrac{8}{25} \right)\left( -\cfrac{5}{12} \right)\implies \cfrac{3}{4}\left( \cfrac{-8}{25} \right)\left( \cfrac{-5}{12} \right) \\ \quad \\ \cfrac{3\cdot (-8)\cdot (-5)}{2\cdot 25\cdot 12}\implies \cfrac{\cancel{3}\cdot (-2\cdot 2\cdot \cancel{2})\cdot (-1\cdot \cancel{5})}{\cancel{2}\cdot (\cancel{5}\cdot 5)\cdot (\cancel{3}\cdot 4)}\)

OpenStudy (jdoe0001):

missed one there... sec

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{4}\left( -\cfrac{8}{25} \right)\left( -\cfrac{5}{12} \right)\implies \cfrac{3}{4}\left( \cfrac{-8}{25} \right)\left( \cfrac{-5}{12} \right) \\ \quad \\ \cfrac{3\cdot (-8)\cdot (-5)}{2\cdot 25\cdot 12}\implies \cfrac{\cancel{3}\cdot (- 2\cdot \cancel{2}\cdot \cancel{2})\cdot (-1\cdot \cancel{5})}{\cancel{2}\cdot (\cancel{5}\cdot 5)\cdot (\cancel{3}\cdot \cancel{2}\cdot \cancel{2})}\)

OpenStudy (jdoe0001):

notice 5* 5 = 25 2*2*2= 8 2*2 = 4

OpenStudy (jdoe0001):

and... ahemm 3*4 or 3*2*2 = 12

OpenStudy (anonymous):

it should come out as 

OpenStudy (anonymous):

\[\frac{ 1 }{ 10}\]

OpenStudy (anonymous):

as the answer

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

well.... I missed one again :( at the botom, a 2... lemme do the whole thing

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{4}\left( -\cfrac{8}{25} \right)\left( -\cfrac{5}{12} \right)\implies \cfrac{3}{4}\left( \cfrac{-8}{25} \right)\left( \cfrac{-5}{12} \right) \\ \quad \\ \cfrac{3\cdot (-8)\cdot (-5)}{2\cdot 25\cdot 12}\implies \cfrac{3\cdot (-2\cdot 2\cdot 2)\cdot ( 5)}{2\cdot (5\cdot 5)\cdot (3\cdot 4)} \\ \quad \\ \cfrac{\cancel{3}\cdot (-1\cdot \cancel{2}\cdot \cancel{2}\cdot \cancel{2})\cdot (-1\cdot \cancel{5})}{\cancel{2}\cdot (\cancel{5}\cdot 5)\cdot (\cancel{3}\cdot \cancel{2}\cdot \cancel{2})}\implies \cfrac{-1\cdot -1}{5} \implies \cfrac{1}{5}\)

OpenStudy (jhannybean):

\[\left(\frac{\color{orange}3}{\color{blue}{4}}\right)\left(\frac{\color{blue}{-8}}{\color{red}{25}}\right)\left(\frac{\color{red}{-5}}{\color{orange}{12}}\right)\] \[\left(\frac{1\cancel{\color{orange}3}}{1\cancel{\color{blue}{4}}}\right)\left(\frac{-2\cancel{\color{blue}{-8}}}{-5\cancel{\color{red}{25}}}\right)\left(\frac{1\cancel{\color{red}{-5}}}{4\cancel{\color{orange}{12}}}\right)\] \[=\frac{-2}{-20} \] \[=\frac{1}{10}\]